Rio Coffee Shoppe sells two coffee drinks—a regular coffee and a latte. The two drinks have the following prices and cost characteristics:

Regular Coffee Latte
Sales price (per cup) $ 1.80 $ 2.80
Variable costs (per cup) 0.60 1.70


The monthly fixed costs at Rio are $8,211. Based on experience, the manager at Rio knows that the store sells 90 percent regular coffee and 10 percent lattes.


Required:

How many cups of regular coffee and lattes must Rio sell every month to break even?

Break Even Sales in Units

Regular Coffee ?

Lattes ?
Answer #1

Contribution margin=Sales-Variable cost

Contribution margin for:

Regular Coffee=(1.8-0.6)=$1.2 per cup

Latte=(2.8-1.7)=$1.1 per cup

Weighted average Contribution margin=Respective Contribution margin*Respective sales mix

=(1.2*0.9)+(1.1*0.1)=$1.19 per cup

Hence total breakeven=Fixed cost/Weighted average Contribution margin

=8211/1.19=6900 cups

Breakeven for:

Regular Coffee=(6900*90%)=6210 cups

Lattes=(6900*10%)=690 cups